2. Consider the following farmer in 2015. She produced an output of $10,000 and consumed 40% of her own output. The rest was sold in the market. She bought $5,000 of consumer goods from the market for her own consumption. She saved whatever money was left in a savings account at a negligible rate of interest ( 1 0%). a. What was the value of her total consumption in 2015? b. What was the value of her 2015 output gap? c. What were the values of her 2015 exports, imports, and current account balance?e d. How much money did she put into her savings account in 2015 e. What was the value of the (i) credit item, and (ii) debit item in her capital account in 2015? f. What was the value of her capital account balance in 2015?
Answer #1

a) Total value of consumption of Farmer = $10000*40%+ $5000= $9000

b) Total output gap value of Farmer = $10000*40%= $4000

c) The value of exports by Farmer = $10000*60%= $6000

The value of imports by Farmer = $5000

Current Account balance of farmer = export value of her output - Import value = $6000-$5000=$1000

d) The amount kept by farmer in her saving account = $6000- $5000= $1000

e) The value of credit item in farmer's capital account = $6000(1+.10) = $6600

The value of debit item in farmer's capital account = $5000

f) The value of capital account balance of farmer = Credit balance -Debit balance = $6600-$5000= $1600